Sara is mixing together a fruit punch for a party. She's made 6 gallons of punch with a mixture of 50% juice. Her mother tells her to change it to a mixture of 60% juice.
How much fruit juice should be added to make the mixture 60% fruit juice (round to the nearest hundredth)?
A)0.4 gallons
B)0.6 gallons
C)1.5 gallons
D)2.1 gallons

Answer by Earlsdon(6294)

Let x = the required number of gallons of pure juce (100% juice) to be added to the 6 gallons of 50% juice mixture.
0.5(6)+x = 0.6(x+6) Simplify.
3+x = 0.6x+3.6 Subtract 0.6x from both sides.
3+0.4x = 3.6 Subtract 3 from both sides.
0.4x = 0.6 Finally, divide both sides by 0.4
x = 1.5 gallons.